# Scattering Cross Section Formula Derived From Macroscopic Model of Detectors

### v1

@@ 1 Introduction
The formula (1), which agrees with empirical data (for further discussion see [4, 8, 41]), is usually theoretically justified by arguing that $\int_{B_{R}}|\Psi_{t}(\boldsymbol{x})|^{2}\,d^{3}\boldsymbol{x}$ (where $B_{R}=\{\boldsymbol{x}\in\mathbb{R}^{3}:|\boldsymbol{x}|<R\}$ denotes the ball of radius $R$ around the origin) should be the probability that no detection has occurred by time $t>0$, and that this quantity changes in $[t,t+dt]$ (to leading order in $R$) by $dt\int_{S_{R}}\sigma(\boldsymbol{x},t)\,d^{2}\boldsymbol{x}$ as given by (1). As we will discuss in Section 2, this justification has some merits but is not satisfactory. Our goal in this paper is to provide a better justification of (1) by deriving that it is, not just a desirable probability density, but the actual probability density of when and where detectors click. Our derivation starts from a model of detectors that is generally applicable, i.e., also without the assumptions of spherical shape and large $R$. We present derivations from two such models, the first using imaginary potentials and the second repeated (“stroboscopic”) nearly-projective position measurements outside $B_{R}$ as effective macroscopic (and approximate) descriptions of detectors. It would be of interest to have a derivation starting from a microscopic model of the detectors as quantum-mechanical many-body systems, but this will not be attempted here; however, an outline of such a derivation can be obtained by combining the reasoning in this paper with the outlined derivation in [40] of imaginary potentials from a microscopic detector model.
@@
The comparison of $T_{D},T_{WID}$, and $T_{WOD}$ is based on a comparison of the evolution of the wave function $\Psi_{WID}$ in our models with the free evolution (i.e., with $V=0$) of the wave function $\Psi_{WOD}$ without detectors for large times $t=\mathcal{O}(R)$ and distances $|\boldsymbol{x}|=\mathcal{O}(R)$, which takes the following well-known asymptotic form, as derived in, e.g., [4, Prop. 3.17], [31, Thm. IX.31], [14, (9.20)]:
@@
A basic difficulty about deriving the scattering cross section (1) arises from the question what to derive it from: one first needs to have a description of detectors that is generally applicable. And this is controversial: not only are there many inequivalent proposals for how to compute the probability distribution of the detection time (e.g., [2, 3, 26]), but also even arguments suggesting the impossibility of such a distribution, such as Pauli’s [29, p. 140] argument that there cannot be any self-adjoint time operator (canonically conjugate to the Hamiltonian) and the quantum Zeno effect, which was formulated by some researchers as leading to the absurd conclusion that “a watched pot never boils” [25]. Our answer to the question about the starting point of the derivation is to use two models of detection (imaginary potentials and repeated position measurements) that are contradiction-free and lead, as we show, to a non-trivial distribution in agreement with (1). $\diamond$
@@
Another difficulty arises because one might expect that the detectors should be modeled as a detecting surface at $S_{R}$, that is, as hard detectors (ones that detect the particle immediately upon arrival, as opposed to soft detectors which may take a while to notice the particle). Hard detectors lead not only to the aforementioned problems of controversial, inequivalent proposals and paradoxes like the quantum Zeno effect, but also, even in paradox-free models of hard detectors such as the absorbing boundary rule [36], to reflection of part of the wave function from the detector surface, which is undesirable and would not fit together with (1). Our answer to this difficulty is to allow soft (but not too soft) detectors for our derivation, as suggested already in [36, Sec. 5]; we have in mind the following reasons:
@@
where $\sigma(\boldsymbol{x}_{1},t_{1},\ldots,\boldsymbol{x}_{N},t_{N})\,d^{2}\boldsymbol{x}_{1}dt_{1}\cdots d^{2}\boldsymbol{x}_{N}dt_{N}$ means the probability that each particle $i=1,\ldots,N$ will be detected in the surface element $d^{2}\boldsymbol{x}_{i}$ and time interval $dt_{i}$ (related formulas are given in [13, Eq.s (33) and (35)]). This formula will also be justified in Section 6. $\diamond$
@@ 2.1 Common Justifications
which is just (1) integrated over $\boldsymbol{x}\in S_{R}$. A similar joint consideration of $\boldsymbol{X}_{D}$ and $T_{D}$ leads to (1). A reasoning of this kind is given in, e.g., [28, Sec. 6.7].
While the result (1) is correct, we argue that the reasoning is not. A first way of seeing this is to notice that the largeness of $R$ and the spherical shape of the detector surface played a role in going from (15) to (16), but not in the derivation of (15); thus, if that derivation was valid, it should apply to arbitrary surfaces instead of large spheres. But for arbitrary surfaces, the right-hand side of (15) may be negative, so (15) cannot be generally true. Moreover, it was shown in [42] that even if the right-hand side of (15) is positive, it is in general not given by a POVM and thus cannot be the distribution of the outcome of any experiment.
@@ 3.2 Second Model: Zeno Dynamics
Figure 2: A plot illustrating the shape of an error function, $\operatorname{erf}(x)$, as defined in Eq. (35).
@@
Our results apply to any limit in which $R\to\infty$, $\mathscr{T}\rightarrow\infty$, $\mathscr{T}/R\rightarrow 0$, $R\sigma_{1}/\mathscr{T}\to\infty$, and $R\sigma_{1}/\mathscr{T}^{2}\to 0$. First, we obtain that (5) holds again. Second, the deviations associated with the Bohmian arrival variables are, with probability approaching 1,
@@ 5.2 Second Approximation
By a Galilean transformation with relative velocity $\hbar k_{0}/m$ [39, Exercise 1.18],
@@
Detection time and place- Essentially, our calculation of the time evolution of $\hat{P}_{\sigma_{n}}\Psi$ justifies the first approximation described in Section 5.1.
@@
In more detail, let us compute the probability of detection at the $n^{\mathrm{th}}$ attempt. We have seen that between $n\mathscr{T}$ and $(n+1)\mathscr{T}$, the step moves outward by $R/n$ and reaches width $w_{n}(\mathscr{T})\ll R/n$; thus, $\Psi$ leaks out of $\Omega$ at most into a shell of thickness $2R/n$; except for the first few values of $n$ (for which $\Psi$ is still concentrated near the origin), this thickness is small compared to $R$ (like $2/n$), so the normalizing constant $\mathcal{N}$ is dominated by the amount of $\Psi$ still inside $\Omega$, or $\mathcal{N}\approx\mathcal{N}_{n}$ as in (79). In particular, the probability of no detection up to $n\mathscr{T}$ is given by $\mathcal{N}_{n}^{-2}$, which means that $T_{D}/R$ is distributed as in (5). Since the thickness is small compared to $R$, $|\boldsymbol{X}_{D}|/R\approx 1$ with high probability, see Figure 4(c). Since the step function acts only on the radial coordinate, the angle dependence of $\Psi$ is unchanged and (5) also applies to $\boldsymbol{X}_{D}/|\boldsymbol{X}_{D}|$; thus, (5) applies completely.
@@
Alternatively, since we already know that the asymptotics of $(T_{WOD},\boldsymbol{X}_{WOD})$ is given by (1) in the sense of (5), in order to derive the same asymptotics for $(T_{D},\boldsymbol{X}_{D})$, it suffices to show that the difference is negligible compared to $R$. To this end and to verify (42)–(45), let us estimate how small the change on $\Psi$ inside $\Omega$ due to the collapse is: Since $|1-\operatorname{erf}(x)|<\exp(-x^{2})$ for sufficiently large real $x$ and $|\operatorname{erf}(e^{i\theta}x)-\operatorname{erf}(x)|<x\exp(-(\text{const.})x^{2})$ as shown after (89), $|1-\operatorname{erf}(e^{i\theta}x)|<\exp(-\alpha x^{2})$ with some $\alpha>0$ for sufficiently large $x$. Thus by time $(n+1)\mathscr{T}$- (i.e., $\Delta t=\mathscr{T}$-), since the step has moved outward by $R/n$ and grown in width to $w_{n}(\mathscr{T})$, the square bracket in (91) is absolutely smaller for $|\boldsymbol{x}|<R$ than $\exp(-\alpha(R/nw_{n}(\mathscr{T}))^{2})$. Since $w_{n}(\mathscr{T})<n\sigma_{1}+\hbar\mathscr{T}/2mn\sigma_{1}<2\max\{n\sigma_{1},\hbar\mathscr{T}/2mn\sigma_{1}\}$ and $n=\mathcal{O}(R/\mathscr{T})$, we have that $R/nw_{n}(\mathscr{T})>\min\{\mathcal{O}(\mathscr{T}^{2}/\sigma_{1}R),\mathcal{O}(\sigma_{1}R/\mathscr{T})\}\gg\mathcal{O}(1)$ by (38), so the square bracket is $\ll\mathcal{O}(1)$, and the perturbation inside $\Omega$ is negligible, which means it can be ignored after the next collapse. Still, at time $n\mathscr{T}+$ for (say) $R-\sigma_{n}\leq|\boldsymbol{x}|\leq R$, the perturbation of $\Psi$ is not negligible, with two consequences: $T_{D}$ might be less than $T_{WID}$ by up to $\mathscr{T}$ (i.e., the particle could be detected before it has even reached $S_{R}$), and $T_{WID}$ might be less than $T_{WOD}$ (as if the particle felt tailwinds): after all, the growth of $w_{n}(\Delta t)$ entails that probability will be transported outward faster than at speed $|\boldsymbol{v}_{WOD}|=\hbar|\boldsymbol{k}_{0}|/m$, at most by the distance $\mathcal{O}(w_{n}(\mathscr{T}))$; since velocities are of order 1 (neither very large nor very small), the particle might arrive earlier by at most $\mathcal{O}(w_{n}(\mathscr{T}))=o(R/n)=o(\mathscr{T})$. In addition, the mere fact that after the particle crosses the sphere $S_{R}$, the next attempted detection may still occur up to $\mathscr{T}$ later can lead to a difference between $T_{D}$ and $T_{WID}$ of order $\mathscr{T}$. We thus obtain (42), (43) (since velocities are bounded and $\boldsymbol{X}_{D}=\boldsymbol{Q}(T_{D})$), (44), and (again since velocities are bounded) (45). As a consequence, we obtain (5).

### v2

@@ 1 Introduction
The formula (1), which agrees with empirical data (for further discussion see [4, 8, 44]), is usually theoretically justified by arguing that $\int_{B_{R}}|\Psi_{t}(\boldsymbol{x})|^{2}\,d^{3}\boldsymbol{x}$ (where $B_{R}=\{\boldsymbol{x}\in\mathbb{R}^{3}:|\boldsymbol{x}|<R\}$ denotes the ball of radius $R$ around the origin) should be the probability that no detection has occurred by time $t>0$, and that this quantity changes in $[t,t+dt]$ (to leading order in $R$) by $dt\int_{S_{R}}\sigma(\boldsymbol{x},t)\,d^{2}\boldsymbol{x}$ as given by (1). As we will discuss in Section 2, this justification has some merits but is not satisfactory. Our goal in this paper is to provide a better justification of (1) by deriving that it is, not just a desirable probability density, but the actual probability density of when and where detectors click. Our derivation starts from a model of detectors. We present derivations from two such models, the first using imaginary potentials and the second repeated (“stroboscopic”) nearly-projective position measurements outside $B_{R}$ as effective macroscopic (and approximate) descriptions of detectors. It would be of interest to have a derivation starting from a microscopic model of the detectors as quantum-mechanical many-body systems, but this will not be attempted here; however, an outline of such a derivation can be obtained by combining the reasoning in this paper with the outlined derivations in [19, 43] of imaginary potentials from microscopic detector models.
@@
The comparison of $T_{D},T_{WID}$, and $T_{WOD}$ is based on a comparison of the evolution of the wave function $\Psi_{WID}$ in our models with the free evolution (i.e., with $V=0$) of the wave function $\Psi_{WOD}$ without detectors for large times $t=\mathcal{O}(R)$ and distances $|\boldsymbol{x}|=\mathcal{O}(R)$, which takes the following well-known asymptotic form, as derived in, e.g., [4, Prop. 3.17], [33, Thm. IX.31], [15, (9.20)]:
@@
A basic difficulty about deriving the scattering cross section (1) arises from the question what to derive it from: one first needs to have a description of detectors that is generally applicable. And this is controversial: not only are there many inequivalent proposals for how to compute the probability distribution of the detection time (e.g., [2, 3, 28]), but also even arguments suggesting the impossibility of such a distribution, such as (i) Pauli’s [31, p. 140] argument that there cannot be any self-adjoint time operator (canonically conjugate to the Hamiltonian) and (ii) the quantum Zeno effect, which was formulated by some researchers as leading to the absurd conclusion that “a watched pot never boils” [27]. Our answer to the question about the starting point of the derivation is to use two models of detection (imaginary potentials and repeated position measurements) that are contradiction-free and lead, as we show, to a non-trivial distribution in agreement with (1). $\diamond$
@@
Another difficulty arises because one might expect that the detectors should be modeled as a detecting surface at $S_{R}$, that is, as hard detectors (ones that detect the particle immediately upon arrival, as opposed to soft detectors which may take a while to notice the particle). Hard detectors lead not only to the aforementioned problems of controversial, inequivalent proposals and paradoxes like the quantum Zeno effect, but also, even in paradox-free models of hard detectors such as the absorbing boundary rule [39], to reflection of part of the wave function from the detector surface, which is undesirable and would not fit together with (1). Our answer to this difficulty is to allow soft (but not too soft) detectors for our derivation, as suggested already in [39, Sec. 5]; we have in mind the following reasons:
@@
where $\sigma(\boldsymbol{x}_{1},t_{1},\ldots,\boldsymbol{x}_{N},t_{N})\,d^{2}\boldsymbol{x}_{1}dt_{1}\cdots d^{2}\boldsymbol{x}_{N}dt_{N}$ means the probability that each particle $i=1,\ldots,N$ will be detected in the surface element $d^{2}\boldsymbol{x}_{i}$ and time interval $dt_{i}$ (related formulas are given in [14, Eq.s (33) and (35)]). This formula will also be justified in Section 6. $\diamond$
@@ 2.1 Common Justifications
which is just (1) integrated over $\boldsymbol{x}\in S_{R}$. A similar joint consideration of $\boldsymbol{X}_{D}$ and $T_{D}$ leads to (1). A reasoning of this kind is given in, e.g., [30, Sec. 6.7].
While the result (1) is correct, we argue that the reasoning is not. A first way of seeing this is to notice that the largeness of $R$ and the spherical shape of the detector surface played a role in going from (15) to (16), but not in the derivation of (15); thus, if that derivation was valid, it should apply to arbitrary surfaces instead of large spheres. But for arbitrary surfaces, the right-hand side of (15) may be negative, so (15) cannot be generally true. Moreover, it was shown in [45] that even if the right-hand side of (15) is positive, it is in general not given by a POVM and thus cannot be the distribution of the outcome of any experiment [11], [42, Sec. 5.1.2].
@@ 3.2 Second Model: Zeno Dynamics
Figure 2: Shape of the error function, $\operatorname{erf}(x)$, as defined in Eq. (35).
@@
Our results apply to any limit in which $R\to\infty$, $\mathscr{T}\rightarrow\infty$, $\mathscr{T}/R\rightarrow 0$, $R\sigma_{1}/\mathscr{T}\to\infty$, and $R\sigma_{1}/\mathscr{T}^{2}\to 0$. The first result is that (5) holds again.
The second result is concerned with the magnitude of the deviations $T_{D}-T_{WID}$ etc.; since these quantities involve the Bohmian trajectories $\boldsymbol{Q}(t)$, defined in (24) and (25), we need to specify how the nearly-projective measurements act for a given Bohmian position $\boldsymbol{Q}(t)$. Here again, we use the notation $\boldsymbol{Q}(t)=\boldsymbol{Q}_{WOD}(t)$ for $\Psi=\Psi_{WOD}$ evolving under the free Schrödinger equation, and $\boldsymbol{Q}(t)=\boldsymbol{Q}_{WID}(t)$ for $\Psi=\Psi_{WID}$ evolving in accordance with the repeated nearly-projective measurements of the observable $1_{\Omega}$. We model the latter as follows, similarly to [38]. At each time $n\mathscr{T}$ of attempted detection, a real-valued random variable $Z$ is chosen independently of the past (including independently of $\boldsymbol{Q}_{WID}(n\mathscr{T}-)$) with probability density
$p_{\sigma_{n}}(z)=\frac{1}{2\sqrt{\pi}\sigma_{n}}\left(1+\operatorname{erf}\left(\frac{z}{2\sigma_{n}}\right)\right)\exp\left(-\frac{z^{2}}{4\sigma^{2}_{n}}\right)\,.$ (42)
The quantity $Z$ serves for representing the fuzziness of the measurement. We define that the outcome of the experiment is
“no detection” if $|\boldsymbol{Q}_{WID}(n\mathscr{T}-)|+Z<R\,$, (43)
“detection” if $|\boldsymbol{Q}_{WID}(n\mathscr{T}-)|+Z\geq R\,.$ (44)
We show in Section 5.2 that this definition leads to exactly the probability $\langle\Psi_{n\mathscr{T}-}|\hat{P}_{\sigma_{n}}|\Psi_{n\mathscr{T}-}\rangle$ for “no detection,” as it should.
The Bohmian position is left unchanged during the collapse, and from time $n\mathscr{T}+$ onwards the Bohmian position moves according to (24), with $\Psi=\Psi_{n\mathscr{T}+}$ given by (41) or (39). We show in Section 5.2 that, conditioned on either “no detection” or “detection” (whichever applies), $\boldsymbol{Q}_{WID}(n\mathscr{T}+)=\boldsymbol{Q}_{WID}(n\mathscr{T}-)$ has distribution density $|\Psi_{n\mathscr{T}+}|^{2}$. This completes the definition of the model of detection. [^3] At time $n\mathscr{T}$, the Bohmian velocity will not have a discontinuity either, as $\Psi_{n\mathscr{T}-}$ gets multiplied by a real valued function (33) or (34), leading to no additional contribution to the total phase. While discontinuities may occur in higher-order time derivatives of $\boldsymbol{Q}(t)$, these do not affect the deviations in the Bohmian variables beyond the errors listed in (45)–(48).
We find that the deviations associated with the Bohmian arrival variables are, with probability approaching 1,
@@ 5.2 Second Approximation
By a Galilean transformation with relative velocity $\hbar k_{0}/m$ [42, Exercise 1.18],
@@
Detection time and place- Essentially, our calculation of the time evolution of $\hat{P}_{\sigma_{n}}^{1/2}\Psi$ justifies the first approximation described in Section 5.1.
@@
In detail, let us compute the probability of detection at the $n^{\mathrm{th}}$ attempt. We have seen that between $n\mathscr{T}$ and $(n+1)\mathscr{T}$, the step moves outward by $R/n$ and reaches width $w_{n}(\mathscr{T})\ll R/n$; thus, $\Psi$ leaks out of $\Omega$ at most into a shell of thickness $2R/n$; except for the first few values of $n$ (for which $\Psi$ is still concentrated near the origin), this thickness is small compared to $R$ (like $2/n$), so the normalizing constant $\mathcal{N}$ is dominated by the amount of $\Psi$ still inside $\Omega$, or $\mathcal{N}\approx\mathcal{N}_{n}$ as in (82). In particular, the probability of no detection up to $n\mathscr{T}$ is given by $\mathcal{N}_{n}^{-2}$, which means that $T_{D}/R$ is distributed as in (5). Since the thickness is small compared to $R$, $|\boldsymbol{X}_{D}|/R\approx 1$ with high probability, see Figure 4(c). Since the step function acts only on the radial coordinate, the angle dependence of $\Psi$ is unchanged and (5) also applies to $\boldsymbol{X}_{D}/|\boldsymbol{X}_{D}|$; thus, (5) applies completely.
@@
An alternative justification can be provided using a Bohmian analysis of the model. Since we already know that the asymptotics of $(T_{WOD},\boldsymbol{X}_{WOD})$ is given by (1) in the sense of (5), in order to derive the same asymptotics for $(T_{D},\boldsymbol{X}_{D})$, it suffices to show that the difference is negligible compared to $R$. To this end and to verify (45)–(48), let us estimate how small the change on $\Psi$ inside $\Omega$ due to the collapse is: Since $|1-\operatorname{erf}(x)|<\exp(-x^{2})$ for sufficiently large real $x$ and $|\operatorname{erf}(e^{i\theta}x)-\operatorname{erf}(x)|<x\exp(-(\text{const.})x^{2})$ as shown after (92), $|1-\operatorname{erf}(e^{i\theta}x)|<\exp(-\alpha x^{2})$ with some $\alpha>0$ for sufficiently large $x$. Thus by time $(n+1)\mathscr{T}$- (i.e., $\Delta t=\mathscr{T}$-), since the step has moved outward by $R/n$ and grown in width to $w_{n}(\mathscr{T})$, the square bracket in (94) is absolutely smaller for $|\boldsymbol{x}|<R$ than $\exp(-\alpha(R/nw_{n}(\mathscr{T}))^{2})$. Since $w_{n}(\mathscr{T})<n\sigma_{1}+\hbar\mathscr{T}/2mn\sigma_{1}<2\max\{n\sigma_{1},\hbar\mathscr{T}/2mn\sigma_{1}\}$ and $n=\mathcal{O}(R/\mathscr{T})$, we have that $R/nw_{n}(\mathscr{T})>\min\{\mathcal{O}(\mathscr{T}^{2}/\sigma_{1}R),\mathcal{O}(\sigma_{1}R/\mathscr{T})\}\gg\mathcal{O}(1)$ by (38), so the square bracket is $\ll\mathcal{O}(1)$, and the perturbation inside $\Omega$ is negligible, which means it can be ignored after the next collapse. Still, at time $n\mathscr{T}+$ for (say) $R-\sigma_{n}\leq|\boldsymbol{x}|\leq R$, the perturbation of $\Psi$ is not negligible, with two consequences: $T_{D}$ might be less than $T_{WID}$ by up to $\mathscr{T}$ (i.e., the particle could be detected before it has even reached $S_{R}$), and $T_{WID}$ might be less than $T_{WOD}$ (as if the particle felt tailwinds): after all, the growth of $w_{n}(\Delta t)$ entails that probability will be transported outward faster than at speed $|\boldsymbol{v}_{WOD}|=\hbar|\boldsymbol{k}_{0}|/m$, at most by the distance $\mathcal{O}(w_{n}(\mathscr{T}))$; since velocities are of order 1 (neither very large nor very small), the particle might arrive earlier by at most $\mathcal{O}(w_{n}(\mathscr{T}))=o(R/n)=o(\mathscr{T})$. In addition, the mere fact that after the particle crosses the sphere $S_{R}$, the next attempted detection may still occur up to $\mathscr{T}$ later can lead to a difference between $T_{D}$ and $T_{WID}$ of order $\mathscr{T}$. We thus obtain (45), (46) (since velocities are bounded and $\boldsymbol{X}_{D}=\boldsymbol{Q}(T_{D})$), (47), and (again since velocities are bounded) (48). As a consequence, we obtain (5).
We still need to verify two properties we asserted in Section 3.2: first, that the probability of “no detection” at $n\mathscr{T}$ agrees with $\langle\Psi_{n\mathscr{T}-}|\hat{P}_{\sigma_{n}}|\Psi_{n\mathscr{T}-}\rangle$; and second the equivariance (preservation of the $|\Psi|^{2}$ distribution) through the collapse. Here is the calculation for the first property. Consider a measurement at time $n\mathscr{T}$. From (42) and the fact that the choice of $Z$ is independent of $\boldsymbol{Q}_{WID}(n\mathscr{T}-)$, we have that
$\mathbb{P}\Bigl(\text{``no detection''}\Big|\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q}\Bigr)$ $=\mathbb{P}\Bigl(|\boldsymbol{Q}_{WID}(n\mathscr{T}-)|+Z<R\ \Big|\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q}\Bigr)$ (97)
$=\mathbb{P}\bigl(Z<R-|\boldsymbol{q}|\bigr)$ (98)
$=\int_{-\infty}^{R-|\boldsymbol{q}|}dz\ p_{\sigma_{n}}(z)$ (99)
$=\frac{1}{4}\left(1-\text{erf}\left(\frac{|\boldsymbol{q}|-R}{2\sigma_{n}}\right)\right)^{2}.$ (100)
Therefore,
$\mathbb{P}(\text{``no detection''})$ $=\int d^{3}\boldsymbol{q}\ \mathbb{P}\Bigl(\text{``no detection''}\Big|\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q}\Bigr)\,\mathbb{P}\bigl(\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q}\bigr)$ (101)
$=\frac{1}{4}\int d^{3}\boldsymbol{q}\ \left(1-\text{erf}\left(\frac{|\boldsymbol{q}|-R}{2\sigma_{n}}\right)\right)^{2}|\Psi_{n\mathscr{T}-}(\boldsymbol{q})|^{2},$ (102)
[…]
which is what we wanted to show.
We turn to the second property. While the Bohmian trajectories are explicitly equivariant under the continuous Schrödinger evolution from time $(n-1)\mathscr{T}+$ up to $n\mathscr{T}$-, measurement at time $n\mathscr{T}$ necessitates clarification on whether this property holds immediately after, i.e., at time $n\mathscr{T}+$. Specifically, we must justify that the distribution of the Bohmian position $\boldsymbol{Q}_{WID}(n\mathscr{T}+)$, conditionally on either “no detection” or “detection” at time $n\mathscr{T}$, still agrees with $|\Psi_{n\mathscr{T}^{+}}|^{2}$. Indeed, using the fact that the Bohmian position remains unchanged during the measurement ($\boldsymbol{Q}(n\mathscr{T}+)=\boldsymbol{Q}(n\mathscr{T}-)$) as well as (100) and (103), we obtain that
$\mathbb{P}\Bigl(\boldsymbol{Q}_{WID}(n\mathscr{T}+)=\boldsymbol{q}\Big|\text{``no detection''}\Bigr)=$
$\qquad=\frac{\mathbb{P}\left(\boldsymbol{Q}_{WID}(n\mathscr{T}+)=\boldsymbol{q},\text{``no detection''}\right)}{\mathbb{P}\text{(``no detection'')}}$ (104)
$\qquad=\frac{\mathbb{P}\left(\text{``no detection''}|\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q}\right)\,\mathbb{P}(\boldsymbol{Q}_{WID}(n\mathscr{T}-)=\boldsymbol{q})}{\mathbb{P}(\text{``no detection''})}$ (105)
[…]
A similar calculation can be done for the event of “detection” at time $n\mathscr{T}$, which completes the proof of the second property. Furthermore, since in the event of “detection” we take for granted that the final position measurement finds the actual position, we have also justified (40).